# Crossroads of Continents: Automated Artifact Extraction for Cultural Adaptation with Large Multimodal Models

### v1

@@ Abstract
In this work, we present a comprehensive three-phase study to examine (1) the effectiveness of large multimodal models (LMMs) in recognizing cultural contexts; (2) the accuracy of their representations of diverse cultures; and (3) their ability to adapt content across cultural boundaries. We first introduce Dalle Street, a large-scale dataset generated by DALL-E 3 and validated by humans, containing $9,935$ images of $67$ countries and $10$ concept classes. We reveal disparities in cultural understanding at the sub-region level with both open-weight (LLaVA) and closed-source (GPT-4V) models on Dalle Street and other existing benchmarks. Next, we assess models’ deeper culture understanding by an artifact extraction task and identify over $18,000$ artifacts associated with different countries. Finally, we propose a highly composable pipeline, CultureAdapt, to adapt images from culture to culture. Our findings reveal a nuanced picture of the cultural competence of LMMs, highlighting the need to develop culture-aware systems. [^1] Dataset and code are available: https://github.com/iamshnoo/crossroads
@@ 1 Introduction
Figure 1: We introduce a large scale balanced dataset for measuring cultural awareness, a task for extracting implicit cultural artifacts from images and a highly composable pipeline for adapting images to different cultural contexts with fine-grained edits.
@@
Culture is hard to define and has always been so. Kroeber (1952) explored how the word has evolved to gain different meanings in different contexts. Recent efforts in natural language processing research have seen growing interest in understanding how culture influences language models and human behavior, including language, art, and decision-making Hershcovich et al. (2022); Adilazuarda et al. (2024); Liu et al. (2024); Ge et al. (2024). As large multimodal models (LMMs; Liu et al., 2023a) intersect more with human life, the need for them to comprehend and respect cultural nuances is crucial. Research in this area explores model alignment with human values, measures of cultural awareness, and methods for cultural adaptation, where content that may be considered as accurately representing a culture or a country, maybe stereotypically, is modified to represent a different one.
@@
The challenge of assessing the capability of understanding and leveraging cultural knowledge in LMMs, given their multimodal nature – processing both text and image data – is significant. Prior research has primarily investigated LMMs for cultural awareness [^2] We maintain that LMMs do not inherently possess human values but that their outputs may display cultural knowledge. by examining their performance on culture-related tasks such as region classification from images Basu et al. (2023); Yin et al. (2023); Pouget et al. (2024), image-caption matching Liu et al. (2021), and cultural image captioning Cao et al. (2024). However, these tasks may not fully determine whether LMMs are responding to cultural cues encoded within their training data or merely identifying superficial culture-associated features. Furthermore, there is a lack of in-depth analysis of how LMMs utilize cultural knowledge during these tasks and whether they exhibit consistent understanding across various cultural contexts.
@@
To address these gaps, We propose a comprehensive three-phase study to deepen our understanding of cultural awareness in LMMs. First, we develop a large-scale dataset for assessing cultural awareness as measured by the ability of LMMs to recognize and differentiate between cultures, with countries as proxies, in a task setting similar to GeoGuessr Geoguessr (2024). Next, we delve deeper into cultural understanding by introducing an artifact extraction task designed to identify implicit cultural artifacts that LMMs use to distinguish between cultures. Finally, based on the insights gained from the artifact extraction task, we propose a cultural adaptation task combining multiple generative models in an end-to-end pipeline to adapt images from one cultural context to another. This pipeline enhances our understanding of LMMs’ adaptability and promises an effective data augmentation technique to improve cultural awareness in LMMs.
Our main contributions are as follows:
@@
  • We introduce, Dalle Street, a collection of $9,935$ images generated by DALL-E 3, covering $67$ countries and $10$ cultural concept classes. Compared to datasets like Dollar Street Rojas et al. (2022), Dalle Street includes more images from underrepresented geographic regions. We validate the appropriateness of country representations with a human study and will release this dataset publicly.
  • We measure how well both open-weight and closed-source LMMs perform on Dalle Street based on the cultural knowledge encoded in them to understand disparities in performance at the geographic sub-region level for a diverse group of concepts and countries. We include a human study baseline to understand where current models may perform better than humans. We also compare LMM performance on two image datasets collected from different sources.
  • We introduce a task for cultural artifact extraction to identify implicit associations. We further develop a method to filter these effectively to discover interesting associations that frequently co-occur for each country.
  • We propose a highly composable end-to-end pipeline, CultureAdapt, using a chain of LMMs to identify a country for a given image, extract culturally relevant artifacts, ground them with bounding boxes, and use diffusion-based inpainting to adapt to a target culture. We show examples of this pipeline and include a CLIPScore-based metric to validate its performance.
@@ 2 Data
We use three different datasets covering a wide variety of cultural concepts, economic ranges, and different sources: (a) Dalle Street: synthetic, DALL-E 3 (OpenAI, 2024) generated; (b) Dollar Street: natural, collected photographs; and (c) MaRVL: scraped from the internet under native speaker guidance. We include count statistics for each of the datasets in Table C.2. Overall, our study uses roughly $20$k images across $3$ datasets, covering $19$ geographical sub-regions and $67$ countries from the United Nations geoscheme UN (2024). Each image is accompanied by a true label for country.
@@
We consider the same $10$ categories and $19$ geographical regions as in Dollar Street and include $67$ countries. [^3] The four extra countries corresponding to new additions to our annotator pool over the timeline of this project. Given these parameters, we generate high resolution images of size $1024$x1024102410241024 from DALL-E 3 OpenAI (2024) for 2 settings - vivid (for generating hyper-real and dramatic images) and natural (for more realistic images). To do so, we follow a templatic prompting approach (prompt shown in Appendix Figure 9) and then sample at least $10$ images for each combination of country and concept class. However, due to global content policy filters imposed on the API calls and limited budget, we end up with $9,935$ images (as opposed to the expected $13,400$). We will release this generated dataset under CC BY-SA 4.0.
We perform a qualitative study, where we ask annotators to determine the “appropriateness” of the generated image on a Likert scale from $1$ (Strongly Disagree) to $5$ (Strongly Agree), where we define appropriateness as the case where the image shown to the annotator is one of the (possibly stereotypical) representations for the given combination of country and category.
@@
This is a dataset of photos of objects and scenes collected by professional and volunteer photographers which specifically focuses on poor and remote environments, containing approximately $200$ concept classes and covering $63$ countries. We filter it for images that do not contain multiple labels, classes which do not cover images for all regions, and classes with subjective naming. Then we refer to our group of annotators from diverse backgrounds to choose the top $10$ categories by the method of collaborative labeling Chang et al. (2017), where effectively we simplify our selection of object classes by choosing the ones which all annotators universally agree on as being a relevant dimension for testing cultural awareness. The final data subset that we end up with thus includes $63$ countries, $19$ geographical regions, $4,137$ images, $10$ concept classes ( car, cups, mugs and glasses, family snapshots, front door, home, kitchen, plate of food, social drink, wall decoration, and wardrobe). This dataset is available under CC BY-SA 4.0.
@@
The primary task for this dataset involves validation of statements about image pairs for five languages: Indonesian, Mandarin Chinese, Swahili, Tamil, and Turkish. We use the images and assign country and region labels to create our task setting for cultural awareness. [^4] MaRVL language to region mapping in Appendix C.1 This dataset gives us $4,914$ images curated by native speakers of each of these languages, across $5$ distinct geographical regions and it is distributed under CC BY 4.0 license for research purposes.
@@ 3 Cultural Awareness (Task 1)
To measure cultural awareness of LMMs, we compare performance on two existing culturally diverse benchmarks and also on Dalle Street. Later on, we also compare this with the baseline from our human study to understand where LMMs stand at this task in comparison to humans. Overall, we find performance varies across sub-regions but both LLaVA and GPT-4V perform nearly equally well.
@@ 3.1 Methods
Given an input image, we prompt two LMMs, LLaVA-NeXT (Liu et al., 2023a), an open-weight model, and GPT-4V vision-preview (OpenAI et al., 2023), a closed-source model, in a zero-shot setting by asking an open-ended question without any possible answer choices: predict the geographical region represented in the image, as per the United Nations geoscheme. We use this geoscheme for three reasons. First, the refusal rate of models is a lot higher when asked about specific country labels. Second, the geoscheme information is already present on English Wikipedia, which is usually included in pre-training corpora for most LLMs. And third, sticking to a geoscheme allows us to parse open-ended generations in a structured format. A potential challenge with this approach is that models may not incorporate the most recent changes in the geoscheme, leading to finer grained errors which would be hard to trace (for example, the Wikipedia page for the geoscheme has been updated more than $20$ times since the cut-off date for the training data of vision-preview, including some changes to island groups like Channel Islands). We focus on countries which have traditionally been consistently classified without changes to geographic region to avoid these inconsistencies.
@@
Table 1: Overall, LLaVA performs as good as or better than GPT-4V on two out of three datasets. Human accuracy on a subset of Dalle Street is $42.63$.
Metric: Accuracy
We process generated text to map it to one of the geographical sub-regions or to a policy violation case, and then compare with true labels by mapping country information to geographical regions, which gives us classification accuracy as a quantitative metric for measuring success. Specifically, we inspect the confusion matrix to better understand on which regions the models make mistakes.
Economic disparities
@@
Human Study - GeoGuessr
We develop an annotation interface similar to Geoguessr (2024) and ask our pool of annotators to label images at either the country or the region level or at continent level. We ask for a maximum of $5$ labels and a minimum of $1$ per image, to adjust for different familiarity levels with multiple randomly sampled regions/cultures. Appendix C.3 includes detailed descriptions of annotator demographics, interface and instructions provided to annotators.
@@ 3.2 Results
Contrary to our expectations, LLaVA is not far off from GPT-4V in terms of cultural awareness, as measured in our experimental settings for this task.
Overall comparison
We show the aggregated results for our Cultural Awareness Task in Table 1. LLaVA performs as good as GPT-4V on Dollar Street and even outperforms it significantly on the Dalle Street images. This is a potential indication that LLaVA might be more prone to forming stereotypical associations between regions and concepts, because the Dalle Street images include such associations to distinguish different geographical regions. As we discuss later, it might be possible to trace this down to the training data.
However, we note that on the MaRVL data, LLaVA performs about as good as random guessing, specifically getting incorrect predictions for nearly all of the images corresponding to Swahili and Turkish. Our hypothesis to explain this is based on the fact that the other two datasets have similar distributions about concepts and countries, whereas MaRVL is much more diverse and covers fewer countries but many indigenous concepts, not all of which the model may have seen during training.
Figure 2: Confusion matrices for GPT-4V on the cultural awareness task for Dalle Street images. Accurate responses match the true subregion exactly. Special labels include Invalid (no match to region labels or incomplete) and ResponsibleAI (no region label due to policy violation). Takeaway: The model shows balanced performance, indicated by a strong leading diagonal. Notably, the model achieves $100$% accuracy for Western Asia images, covering Iran, Jordan, Lebanon, Oman, Palestine, and Turkey.
Subregion level analysis
Figure  2 (and Figure  C.4.15) indicates that both GPT-4V and LLaVA have strong leading diagonals, which is an indicator that they get many predictions correct. GPT-4V has many cases where it does not provide any answer due to violations of the content policy filter. Instead of trying to find workarounds for the filter, we show it as is, because this would represent the experience of any typical user of the system. Very interestingly, both GPT-4V and LLaVA gets all of the answers from Western Asia correct. In general, LLaVA makes a lot of mistakes for different regions, but defaults to South America as the incorrect answer for these cases, whereas for GPT-4V the same translates to the ResponsibleAI policy filter output as the incorrect answer. We include results for the other two datasets in the Appendix (Figure C.4.16 and  C.4.17 ), and find similar trends.
Figure 3: We normalize income data from Dollar Street into region specific quartiles and plot corresponding accuracies for GPT-4V. Takeaway: Lower income quartiles (blue, orange bars) are more accurate in Africa and Asia, but higher income quartiles (red, green) are more accurate in Americas; in Europe all quartiles have similar accuracies.
Economic disparity - GPT-4V
We utilize the monthly income information available as part of the Dollar Street data, and divide the images into groups based on normalized income quartiles for each broad geographical region - Africa, Asia, Americas, Europe and Oceania (see Figure 3). Lower income quartiles show better performance for Africa and Asia, whereas for America performance improves for higher income groups. This discrepancy reveals the latent biases of the model, which seem to associate Africa with poorer contexts and America with richer ones. For Europe, performance is nearly similar for all quartiles.
4 Automated Extraction of Implicit Cultural Artifacts (Task 2)
To identify the implicit associations that the models may be using for performing Task $1$, we propose a second task to extract specific cultural artifacts from the images, and discover interesting associations, which are usually stereotypical for the relevant countries.
Figure 4: We assign a score for each identified artifact based on how likely it is to co-occur for a given country. Across all such scores, we find the mean and standard deviation to define a range. Scores that lie outside this range would correspond to items that co-occur frequently for a given country, and would most likely represent stereotypical implicit associations.

### v2

@@ Abstract
We present a comprehensive three-phase study to examine ($1$) the cultural understanding of Large Multimodal Models (LMMs) by introducing Dalle Street, a large-scale dataset generated by DALL-E 3 and validated by humans, containing $9{,}935$ images of $67$ countries and $10$ concept classes; ($2$) the underlying implicit and potentially stereotypical cultural associations with a cultural artifact extraction task; and ($3$) an approach to adapt cultural representation in an image based on extracted associations using a modular pipeline, CultureAdapt. We find disparities in cultural understanding at geographic sub-region levels with both open-source (LLaVA) and closed-source (GPT-4V) models on Dalle Street and other existing benchmarks, which we try to understand using over $18{,}000$ artifacts that we identify in association to different countries. Our findings reveal a nuanced picture of the cultural competence of LMMs, highlighting the need to develop culture-aware systems. [^1] Dataset and code are available: https://github.com/iamshnoo/crossroads
@@ 1 Introduction
Figure 1: We introduce a large-scale dataset for measuring cultural awareness, an artifact extraction task for implicit cultural associations, and a modular pipeline for culturally adapting images with fine-grained edits.
@@
Culture is hard to define and has always been so. Kroeber (1952) explored how the word has evolved to gain different meanings in different contexts. Recent efforts in natural language processing research have seen growing interest in understanding how culture influences language models and human behavior, including language, art, and decision-making Hershcovich et al. (2022); Adilazuarda et al. (2024); Liu et al. (2024); Ge et al. (2024). As large multimodal models (LMMs) intersect more with human life, the need for them to comprehend and respect cultural nuances is crucial. Research in this area focuses on model alignment with human values, assessing cultural awareness, and exploring cultural adaptation; the goal is to modify content that represents one culture or country, often stereotypically, to reflect a different one, to suit audiences from different cultural backgrounds better.
@@
The challenge of assessing the capability of understanding and leveraging cultural knowledge in LMMs is significant. Prior research has primarily investigated LMMs for cultural awareness [^2] We maintain that LMMs do not inherently possess human values but that their outputs may display cultural knowledge. by examining their performance on tasks such as region classification from images Basu et al. (2023); Yin et al. (2023); Pouget et al. (2024), image-caption matching Liu et al. (2021), and cultural image captioning Cao et al. (2024). However, these tasks do not determine whether LMMs are responding to cultural cues encoded within their training data or merely identifying superficial cultural associations.
@@
To address these gaps, first, we develop a new large-scale dataset to assess cultural awareness as measured by the ability of LMMs to recognize and differentiate between cultures, with countries as proxies, in a task setting similar to GeoGuessr Geoguessr (2024). Next, we introduce a task designed to identify implicit associations between cultures and artifacts Liu et al. (2024) that LMMs use to distinguish between cultures. Finally, we propose a cultural adaptation framework combining multiple generative models in an end-to-end pipeline to adapt images from one cultural context to another by modifying the underlying implicit associations. Our main contributions are as follows:
@@
  • Dataset: We introduce Dalle Street, a collection of $9{,}935$ images generated by DALL-E 3, covering $67$ countries and $10$ cultural concept classes, with more images from underrepresented geographic regions compared to datasets like Dollar Street Rojas et al. (2022).
  • Benchmark: We measure how well humans and multimodal large language models (open- and closed-source) can identify countries for images in Dalle Street and two other datasets (Dollar Street, MaRVL) to study disparities in performance at the geographic subregion level for a diverse group of concepts and countries.
  • Task: We introduce a task for identifying implicit associations by extracting cultural artifacts from images and filtering them to discover associations that frequently co-occur for each country.
  • Framework: We propose a modular end-to-end pipeline, CultureAdapt (Figure 10), to adapt an image to a target culture by updating identified implicit cultural associations in it using diffusion-based inpainting. We evaluate results by introducing a CLIPScore-based metric.
@@ 2 Data
We study around $20$k images from three datasets (data statistics in Table 2) covering a wide variety of cultural concepts, economic ranges, and data sources: (a) Dalle Street: synthetic, DALL-E 3 OpenAI (2024) generated; (b) Dollar Street: natural, collected photographs; and (c) MaRVL: web-scraped under native speaker guidance. All the datasets are available under CC BY-SA $4.0$ license.
@@
We use $10$ concept classes (car, family snapshots, front door, home, kitchen, plate of food, cups/mugs/glasses, social drink, wall decoration, and wardrobe), $19$ geographical regions, and $67$ countries (Section A.1), similar to Dollar Street. We generate $1024\times 1024$ images with DALL-E 3 OpenAI (2024) in two styles - vivid (hyper-realistic) and natural (realistic), prompting with a template (Figure 14) that specifies the concept class and target country. At least $10$ images are sampled per country-concept combination, yielding $9{,}935$ images after filtering out content policy violations from API calls. A qualitative study on a randomly sampled subset of around $300$ generated images and $14$ participants (Table 1(a)) shows most annotators agree that the images reflect stereotypical country representations, with less than $1\%$ of images receiving strong disagreement. When unsure, participants tend to neither agree nor disagree about the “appropriateness” of an image. Our annotators also mark visual cues in these images, including both explicit (e.g., flags) and implicit (e.g., color schemes) cultural artifacts. This feedback motivates our artifact extraction task and its use in our cultural adaptation framework.
@@
This is a dataset of photos of objects and scenes collected by professional and volunteer photographers. We filter it for images that do not contain multiple labels, classes that do not cover images for all regions, and classes with subjective naming. Then we refer to our group of annotators from diverse backgrounds to choose the top $10$ categories by the method of collaborative labeling Chang et al. (2017), where we simplify our selection of object classes by choosing the ones which all annotators universally agree on as being a relevant dimension for testing cultural awareness. Our data from this source includes $4{,}137$ images from $63$ countries, $19$ geographical regions, across $10$ concept classes.
@@
This is originally a dataset for validation of statements about image pairs curated by native speakers in five languages: Indonesian, Mandarin Chinese, Swahili, Tamil, and Turkish. We assign country and region labels (based on corresponding languages) to get $4{,}914$ images across $5$ geographical regions.
@@ 3 Cultural Awareness (Task 1)
We compare performances of humans, LLaVA and GPT-4V on Dalle Street, Dollar Street and MaRVL in terms of their ability to predict the country given an image. Overall, we find performances vary across sub-regions, but both LLaVA and GPT-4V perform better than humans.
Appropriateness Category Percentage (%)
Agree $40.79$
Neither Agree nor Disagree $34.54$
Strongly Agree $19.41$
Disagree $4.61$
Strongly Disagree $0.66$
(a) Results from our human study on appropriateness for generated images show that most participants agree or are neutral, with less than $1$% expressing strong disagreement.
Geographical Level Accuracy (%)
Country Level $22.16$
Subregion Level $47.63$
Continent Level $77.77$
Union Accuracy $78.03$
Intersection Accuracy $21.91$
(b) Accuracy for the cultural awareness task improves from country to subregion to region level.
Table 1: (a) Perceived appropriateness of generated images by human participants. (b) Cultural awareness accuracy at different geographical levels for a subset of Dalle Street images.
@@ 3.1 Methods
Given an input image, we prompt LLaVA-NeXT Liu et al. (2023a) and GPT-4V vision-preview OpenAI et al. (2023) in a zero-shot generative setting by asking an open-ended question without providing answer choices: Predict the geographical region represented in the image, as per the United Nations geoscheme UN (2024). We use this geoscheme for three reasons: ($1$) models have a higher refusal rate when queried with specific country labels; ($2$) the geoscheme is included in most LLM pre-training data (English Wikipedia); and ($3$) it enables structured parsing of open-ended generations. We focus on countries with stable geographic classifications to prevent errors due to geoscheme updates.
Evaluation Metrics
We process generated text to map it to one of the geographical sub-regions or a policy violation case and then compare it with true labels by mapping country information to geographical regions, which gives us classification accuracy as a quantitative metric for measuring success. Since this is a typical classification problem, we also inspect the confusion matrix to locate sub-regions with more errors.
Economic disparities
For Dollar Street, we also have data available for the monthly income of the family corresponding to each image. We use this information to understand differences in performance across economic groups by looking at region-specific normalized income quartiles.
@@
Figure 2: LLaVA matches or outperforms GPT-4V on two of three datasets. Human accuracy on a Dalle Street subset is 47.63%.
Figure 3: Confusion matrices for GPT-4V on the cultural awareness task for Dalle Street images. Accurate responses match the true subregion. Special labels include Invalid (no match or incomplete) and ResponsibleAI (policy violation). Takeaway: The model performs well, with a strong leading diagonal and $100\%$ accuracy for Western Asia (which covers Iran, Jordan, Lebanon, Oman, Palestine, Turkey).
Human Baseline
$14$ annotators label images at the country, subregion, or continent level, with $1$ to $5$ guesses per image, accounting for varying familiarity with different regions. We first evaluate exact match accuracy at the country, then subregion, and finally continent levels. We also consider two cases: union (the correct answer appears at any level) and intersection (the correct answer appears at all levels). Table 1(b) shows that while country-level accuracy is low, it nearly doubles at each broader geographic level (Table 3).
3.2 Results
We find similar trends across datasets for both models, with some variations across subregions.
Figure 4: We normalize Dollar Street income data into region-specific quartiles and plot accuracies for GPT-4V. Takeaway: Lower income quartiles (Q1111, Q2222) show higher accuracy in Africa and Asia, while higher quartiles (Q3333, Q4444) perform better in the Americas. In Europe, accuracy is similar across all quartiles.
Overall comparison
LLaVA performs as good as GPT-4V on Dollar Street and outperforms it significantly on the Dalle Street images (Table 3). This indicates that LLaVA may have implicitly learned stereotypical associations between regions and concepts because the Dalle Street images include such associations (Section 4). However, on the MaRVL data, LLaVA performs about as good as random guessing. This may be because MaRVL covers specific indigenous concepts that the model may not have seen before.
Subregion Level Analysis
GPT-4V performs well on Dalle Street, with a strong leading diagonal indicating many correct predictions (Figure 3). However, it often provides no answer due to content policy violations. Notably, both models accurately predict all Western Asian images (Figure 23). LLaVA tends to default to South America for incorrect answers, while GPT-4V defaults to policy violations. Similar trends are observed in other datasets (Figures 24, 25).
Economic Disparity
Using income data from Dollar Street, we group results by normalized income quartiles across Africa, Asia, Americas, Europe, and Oceania (GPT-4V - Figure 4, LLaVA - Figure 22). Performance is better for lower-income quartiles in Africa and Asia, while it improves with higher-income groups for the Americas. This might indicate that the model defaults to associating Africa with poorer contexts and America with wealthier ones. For Europe, performance remains consistent across all quartiles.
4 Extracting Implicit Associations of Cultures and Artifacts (Task 2)
We propose to extract cultural artifacts (material items) from the generated images to identify the implicit associations the models may use for Task $1$. We find associations that are usually stereotypical (and not truly representative) for the relevant countries. This provides a better understanding of the models, enabling us to develop our approach for Task $3$ for cultural adaptation of images.
Figure 5: We score each artifact based on its likelihood of co-occurrence for a country. Scores outside the mean and standard deviation range (red) indicate frequent co-occurrences, representing implicit (potentially stereotypical) associations.
4.1 Methods
We use GPT-4V vision-preview for open vocabulary object detection Zareian et al. (2021), using a detailed prompt (Figure 16) to extract information about concept classes in Dalle Street images, including descriptions, color, [^3] This refers to object and person appearance, not race. and person count. GPT-4V’s strong instruction-following capabilities result in nearly perfect JSON outputs, which we lightly post-process and summarize using GPT-$4$ turbo (prompt in Figure 17). This process yields many unique associations for each country. In our initial experiments, GPT-4V significantly outperformed LLaVA, hence we report GPT-4V results.
Salient associations
To identify salient artifacts that appear more frequently in one country than others (potentially stereotypical associations), we follow an approach similar to  Jha et al. (2024): compute the term frequencies of each artifact for each country and also compute document frequency as the number of times an artifact occurs across all countries, to calculate a tf-idf score by multiplying term frequency and the inverse of the document frequency. We then perform a qualitative evaluation of outliers from the distribution of these scores.
Figure 6: We identify more than $18{,}000$ unique cultural artifacts across all countries as part of our second task and then filter them to find salient ones (Table 4). This figure shows the strongest correlated artifacts for $5$ randomly picked countries. We include more such examples of country-object associations in the Appendix (Figure 26).
Evaluations
@@ 4.1 Methods
Figure 7: Human validation of a random sample of generated artifacts shows (1) low hallucination (high presence scores) and (2) more than a random fraction of artifacts usually display cultural relevance.
Color Associations for Countries
We calculate the mean RGB vector for each Dalle Street image, then average them to get a global mean vector. We repeat this process at the country level and measure each country’s distance from the global mean, identifying colors more strongly associated with specific countries across the three RGB dimensions.
Counting the Number of People
We observed that DALL-E 3 generates varying population densities across countries for identical prompts. To explore this, we use an object detection prompt (Figure 16) to count people in each image, split into three buckets: less than $5$, $5$ to $10$, and more than $10$ people. We process related terms (e.g., people, person, man, woman) and aggregate country-level statistics to analyze population density distributions in the generated images. Annotators validate a random sample of these counts, finding general consistency, though they often do not reflect real-world population densities accurately.
Figure 8: We explore how countries are distributed on a color spectrum by first calculating a global average RGB vector and then defining deltas along each axes aggregated at the country level. Takeaway: We find interesting associations - Greece is strongly correlated with blue, and the Netherlands with red/orange.
@@ 4.2 Results
We analyze all the DALL-E 3 generated images to discover implicit associations between countries and cultural artifacts. Our method effectively surfaces implicit associations but also highlights challenges with stereotype reinforcement and demographic inaccuracies.
Artifact Associations
Our analysis shows that some cultural artifacts are strongly tied to specific countries, with certain artifacts exceeding one standard deviation from the mean tf-idf score (Figure 5). While these associations can offer cultural insights, they often reflect negative stereotypes, such as the over-representation of palm trees for tropical regions, which overlooks broader diversity. This suggests that while our method captures implicit associations, it also underscores the need to refine models to avoid reinforcing such stereotypes.
Color Associations
Models not only associate cultural artifacts with countries but also colors. In Figure 8, the RGB delta values for several countries in Dalle Street fall outside the standard deviation. For instance, Greece is strongly associated with blue and the Netherlands with red, likely due to recurring elements like blue seas in Greek images and red tulip fields in Dutch ones.
People-Count Associations
Most images fall into the extreme buckets (less than $5$ or more than $10$ people), with few in the middle (Figures 28,9), often misrepresenting actual population densities. In general, African countries tend to fall into high person-count buckets, while European ones are on the low person-count end, possibly reflecting the model’s perception of collectivist versus individualistic societies.
5 Cultural Adaptation (Task 3)
We propose a method to edit a given image for a target culture by modifying the detected salient implicit associations between countries and artifacts.